OpenStudy (anonymous):

help me .... inverse Laplace transform of {1}

OpenStudy (amistre64):

um, just integrate e^(-st) dt

OpenStudy (amistre64):

from 0 to inf of course

OpenStudy (amistre64):

\[\int e^{-st}dt=\frac{1}{s}(-e^{-s\infty}+e^{-s0})\]

OpenStudy (anonymous):

it is answer for lapace transfrom not inverse

OpenStudy (amistre64):

really ... hmmm, i guess i need better glasses

OpenStudy (anonymous):

:)

OpenStudy (amistre64):

interesting, the wolf says it something called a dirac delta function

OpenStudy (amistre64):

which seems to be related to the heaviside step function

OpenStudy (amistre64):

i often find pauls site readable http://tutorial.math.lamar.edu/Classes/DE/DiracDeltaFunction.aspx i dont know enough about this topic to offer anything else

OpenStudy (anonymous):

thank u

OpenStudy (amistre64):

have fun with it :)

OpenStudy (anonymous):

i think the slove is ( s )